A section of blood vessel wall has become weakened resulting in an increase in the vessel wall diameter of 32%.

For this question treat blood as an incompressible non viscous fluid of density 1050 kg m−3. If the vessel normally carries blood at a speed of 1.2 m s−1 then what is the magnitude of the increase in pressure in the bulge in the blood vessel? (to 2 s.f and in Pa)
1
Expert's answer
2020-04-22T10:05:16-0400

Let us use Bernoulli's equation and the fact that the flux of the blood is continuous.

The flux of is time "\Delta t" at two points with different diameter is equal, hence "S_1 v_1 \Delta t = S_2 v_2 \Delta t", or in terms of diameter, "\pi \left(\frac{d_1}{2}\right)^2 v_1 \Delta t = \pi \left(\frac{d_2}{2}\right)^2 v_2 \Delta t", from where "v_2 = \left(\frac{d_1}{d_2}\right)^2 v_1", where "v_1" is the speed before increase in diameter.

According to Bernoulli's equation: "p_1 + \rho g h_1 + \frac{\rho v_1^2}{2} = p_2 + \rho g h_2 + \frac{\rho v_2^2}{2}". Since the height in both cases is the same, terms "\rho g h_1, \rho g h_2" cancel and therefore "\Delta p = p_2 - p_1 = \frac{\rho}{2}[v_1^2 - v_2^2] = \frac{\rho v_1^2}{2}\left[1 - \left(\frac{d_1}{d_2} \right)^4\right]" .

New diameter is "d_2 = (1+\eta)d_1", where "\eta = 0.32" (corresponds to 32%).

Hence, finally "\Delta p = \frac{\rho v_1^2}{2}\left[1 - \frac{1}{(1+\eta)^4}\right] \approx 506.9 Pa".